# DrVoice: Parallel Speech-Text Voice Conversation Model via Dual-Resolution Speech Representations

### v2

@@ 3.1 Speech Tokenization and Detokenization
3.2 Multimodal Large Language Model (MLLM)
Parallel Joint Speech-Text Model.
$c_{t}=E_{\text{speech}}(s_{t})+E_{\text{text}}(t_{t})$ (1)
where $E_{\text{speech}}$ and $E_{\text{text}}$ denote the embeddings of speech and text tokens, respectively. Since the lengths of speech tokens and text tokens are typically mismatched, the shorter sequence is padded with special token <|SIL|>to align them for each utterance. The autoregressive generation process is as follows:
$P(y_{t}|y_{<t},x)=\prod_{i=1}^{t}P(y_{i}|y_{<i},x)$ (2)
where $x$ is the input sequence and $y_{t}=(s_{t},t_{t})$ denotes the joint speech-text output at timestep $t$. This unified formulation enables seamless integration of speech and text generation within a single autoregressive framework. To preserve the intrinsic linguistic understanding and generation capabilities of pretrained text-LLMs while enabling cross-modal interactions, three key methodological components are introduced for intermodal coordination.
Inner Text Streaming.
Inspired by Moshi (Défossez et al., 2024), explicit text streaming is incorporated to aid speech generation as a common semantic scaffold. We focus on performing modality alignment exclusively at the assistant end. This design adheres to the asymmetric nature of human-machine interactions: while user inputs are typically unimodal (either text or speech), the assistant’s responses can be a coordinated multimodal output.
Speech Token Grouping.
A crucial parameter for discrete tokens is sampling rate, which determines the input/output sequence length. GLM-4-Voice (Zeng et al., 2025) investigates sampling rates from 6.25Hz to 50Hz on LibriSpeech, revealing minimal differences in Word Error Rate (WER) between 50Hz and 25Hz but significant degradation at 12.5Hz and 6.25Hz. Hence, our work adopts 25Hz sampling rate. To resolve the temporal resolution mismatch between speech signals (25Hz) and text tokenization ($\sim$3Hz) (Chen et al., 2024), a grouping mechanism is designed:
where $\mathbf{s}_{j}$ denotes speech tokens, $\|$ represents feature concatenation, and $k$ is the grouping factor determined by the ratio between speech token rates and text token rates. This design compresses the speech token length from $T$ to $T/k$ and the grouped speech representations align better with text. Different from Chen et al. (2024) that employs linear projection of audio logits into group-sized representations for parallel multi-token prediction, DrVoice strategically designs an ungrouping mechanism and incorporates a dedicated Speech Refined Head (SRH) to enable autoregressive generation of individual speech tokens.
Speech Refined Head.
SRH is proposed to enhance speech generation capabilities, which utilizes the last hidden state of the shared LLM layer (SLLM) as conditional input and incorporates contextual speech information to autoregressively generate speech tokens. While conventional speech grouping strategies–which cluster speech tokens into semantically meaningful units–have proven effective for speech recognition and understanding tasks (Chen et al., 2024; Zhang et al., 2024b), our experiments reveal their inherent limitations in generative scenarios, since speech token grouping inevitably loses some fine-grained acoustic details. To address this limitation, an ungrouping process is performed as follows: the SLLM’s final hidden state is mapped to group-sized embeddings via linear projection
and time splitting
where $\mathbf{h}_{ug}^{(i)}\in\mathbb{R}^{d_{ug}/k}$, the resulting $\mathbf{H}$ serves as conditional input for SRH that autoregressively generates speech tokens. Following our practice on SLLM, representations of preceding speech tokens and $\mathbf{H}$ are aggregated. Speech token prediction is trained to maximize the conditional probability:
The end-to-end training objective integrates the two losses through multi-task learning paradigm:
$\mathcal{L}_{\text{TH}}=-\sum_{i=1}^{T}\log P(t_{i}|c_{<i},\mathbf{g}),$ (8)
where $\mathcal{L}_{\text{TH}}$ is the autoregressive loss over text head, $\lambda$ and $\mu$ are hyperparameters.
3.3 CoM-Mixing Training strategy
The chain-of-modality (CoM) methodology (Zhang et al., 2023a) can enhance performance by leveraging intermediate textual transcriptions, which is particularly suitable for scenarios where real-time processing is not critical. Furthermore, practical applications involve scenarios where both speech and text output or text-only output are required, necessitating the system to dynamically generate different modalities based on specific needs. Seven interaction patterns for diverse input-output requirements are summarized in Table 1. System prompts are employed to regulate the model’s output behavior. For example, prompts such as ‘‘[System] You are a helpful assistant and asked to generate both text and speech tokens at the same time.’’ explicitly guide the model to produce multimodal outputs. Detailed prompts are shown in the Appendix A. During inference, specifying these system prompts enables the generation of desired output results. For CoM-Mixing training, we construct data variants following the seven interaction patterns and obtain a mixture of data for training the model. During inference, user can manually prepend the appropriate system prompt to the input sequence to meet the output requirement. This flexibility ensures adaptability to varying modality generation demands.
@@ 4 Experiments
Table 2: Speech-to-Text (S $\rightarrow$ T) and Speech-to-Speech (S $\rightarrow$ S) generation performance on spokenQA benchmarks. The results for the S2T baselines are taken from the previous SOTA model Kimi-Audio’s paper (Xu et al., 2025), and the S2S results are taken from the previous SOTA model VITA-Audio’s paper (Long et al., 2025). Otherwise, results are taken from their original papers. The model size refers only to the parameter count of the LLM backbone, excluding the audio encoder, audio tokenzier, and audio detokenzier module. Frame Rate (In/Out) denotes the input speech frame rate and the output speech plus text frame rate for the LLM backbone. The average number of text tokens corresponding to one second of speech is denoted by $\tau$. Metrics are benchmark-specific (higher is better). The Avg. score averages two benchmarks. The best results in each group are highlighted in bold, while the second-best results are underlined.
Model Size Frame Rate (In/Out) Llama Q. Web Q. Avg.
$\rightarrowT S\rightarrowS$ $\rightarrowT S\rightarrowS$ $\rightarrowT S\rightarrowS$
@@
DrVoice-Small (Our) 2B 5/5 73.0 56.0 42.1 31.1 57.5 43.5
Moshi (2024) 7B 12.5/12.5 62.3 21.0 26.6 9.2 44.5 15.1
@@
LUCY (S2) (2025) 7B 12.5/48 57.3 51.0 29.3 18.2 43.3 34.6
[…]
Kimi-Audio (2025) 8B 12.5/12.5 79.3 - 70.0 - 74.7 -
[…]
Stream-Omni (2025) 7B 25/25 76.3 65.0 44.2 27.5 60.3 46.3
[…]
DrVoice (Our) 8B 5/5 85.0 71.0 65.2 40.1 75.1 55.5
Table 3: Speech quality performance on the collections of Spoken Quesion Answering Benchmarks following Zeng et al. (2025), in terms of UTMOS for assessing overall speech quality and ASR WER for assessing alignment between generated speech and text. Baseline scores are copied from Zeng et al. (2025). FR(In/Out) indicates the input speech frame rate and the output (speech + text) frame rate for the LLM backbone, while $\tau$ represents the average text tokens per second of speech. The best results in each group are bolded.
[…]
Mini-Omni (2024a) 0.5B 50/48 3.17 25.28
DrVoice-Small 2B 5/5 3.87 8.86
[…]
Moshi (2025) 7B 12.5/12.5 3.90 7.95
[…]
DrVoice 8B 5/5 3.86 6.75
Table 4: Impact of the Grouping Factor on Llama Q. S2T and S2M performance. Experiments were conducted on DrVoice-Small w/o. CSE, using data subsets for faster exploration. The S2T model was trained exclusively on mixed S2T and T2T data, while the S2M model was trained on mixed S2M and T2T data.
Grouping Factor S2T S2M (T/S)
1 55.67 4.00 / 2.67
3 64.67 15.67 / 5.00
5 63.33 37.67 / 28.00
7 62.67 36.00 / 16.67
@@
Figure 2: Computational Resources across different Grouping Factor.
@@ 4.1 Experimental Setup
Datasets.
Approximately 210K hours of audio-text paired data for speech-text modality synchronization is adopted for pre-training speech refined head. 17.8K hours of English speech conversations data synthesized from text are adopted for DrVoice training: (1) VoiceAssistant (Xie and Wu, 2024a, b), which contains human-agent interactions; (2) Spoken Alpaca  [^1] https://huggingface.co/datasets/GSQA/spoken-alpaca-gpt4, which provides instructed dialogues; and (3) Synthesized Datasets, which are synthesized from text datasets such as Infinity-Instruct (BAAI, 2024), using the CosyVoice TTS system (Du et al., 2024a, b). Following the common practice in prior works (Long et al., 2025; KimiTeam et al., 2025; Yao et al., 2024; Zhang et al., 2025), the Speech-to-Speech ($S\rightarrow S$) and Speech-To-Text ($S\rightarrow T$) performance is evaluated on widely used Spoken Question Answering benchmarks, including Web Questions (Berant et al., 2013) and Llama Questions (Nachmani et al., 2024).
Evaluation Metrics.
@@
Accuracy is emplyed to assess model performance on datasets. The generated speech is transcribed to text using whisper-v3-large model (Radford et al., 2022), then word error rate (WER) of transcripts is computed against the generated text to assess the alignment between generated speech and text, as well as using UTMOS (Saeki et al., 2022) to evaluate the overall speech quality, following Zeng et al. (2025).
Baselines.
@@
Representive baselines are selected to cover diverse modeling paradigms: the cascaded system Audio Flamingo 3 (Zhang et al., 2025), Text-Driven models including Freeze-Omni (Défossez et al., 2024), MinMo (Chen et al., 2025), MiniCPM-o 2.6 (Yao et al., 2024), Llama-Omni2 (Fang et al., 2025), and Qwen2.5-Omni (Xu et al., 2025), and among Joint Speech-Text models, interleaved models including Baichuan-Omni-1.5 (Li et al., 2025), GLM-4-Voice (Zeng et al., 2025), Stream-Omni (Zhang et al., 2025), and VITA-Audio (Long et al., 2025), alongside parallel models including LUCY (Gao et al., 2025), Moshi (Défossez et al., 2024), and Kimi-Audio (Xu et al., 2025). Noting that Kimi-Audio achieves state-of-the-art (SOTA) performance on Speech-to-Text Question Answering, while VITA-Audio sets the SOTA on Speech-to-Speech Question Answering. This configuration enables systematic comparisons across mainstream speech-text modeling strategies.
Implementation Details.
Qwen2.5-0.5B is trained following $T2M$ paradigm with Speech-text Alignment data to initialize SRH (SRH-PT). DrVoice uses Qwen2.5-7B-Instruct as its base LLM, while DrVoice-Small utilizes Qwen2.5-1.5B-Instruct as the base LLM. Grouping factor is set to $k=5$. The warmup rate is set to $1\%$ of the total training steps. For DrVoice, the max learning rate is set to $1\times 10^{-5}$ and then decayed via cosine annealing to $1\times 10^{-6}$, while for DrVoice-Small and SRH-PT it is set to $1\times 10^{-4}$ and then decayed to $1\times 10^{-5}$. The AdamW method (Loshchilov and Hutter, 2019) is used for optimization. Experiments are run on 32x NVIDIA Tesla A800 80G GPUs with Brain floating-point format (BF16(Kalamkar et al., 2019) to accelerate training and decoding. For DrVoice, DeepSpeed ZeRO-2 (Rajbhandari et al., 2020) is implemented to prevent GPU out-of-memory. It takes about $20$ hours on SRH-PT, and about $45$ hours on DrVoice post training.
4.2 Results and Discussions
@@
The quality of generated speech is evaluated in terms of overall speech quality (UTMOS) and alignment between generated speech and text (ASR-WER(%)), as shown in Table 3. Among models with comparable sizes, DrVoice achieves low ASR-WER, indicating strong alignment when generating both speech and text simultaneously. In contrast, UTMOS scores are more constrained by the speech tokenizer and detokenizer (Hassid et al., 2023; Du et al., 2024a). The gap in UTMOS between DrVoice and GLM-4-Voice suggests room for improvement in speech naturalness and intelligibility.
We conduct extensive ablation study to analyze effects of critical algorithmic designs.
Continuous Speech Encoder.
@@
As shown in Table 5, the Continuous Speech Encoder (CSE) is vital for tasks involving speech inputs. Removing it from DrVoice-Small leads to a significant performance drop in speech understanding and speech generation. Specifically, the S2T score decreases from 72.33 to 62.33 (13.8% relatively), and the S2M (T) score falls from 68.67 to 61.67 (10.2% relatively). In contrast, the impact on text-only tasks is minimal, with the T2T score only slightly decreasing from 75.33 to 74.00. This confirms the CSE’s effectiveness in processing and representing speech information.
Dual-Resolution.
Our dual-resolution approach combines input grouping for efficiency and comprehension with a refinement head for high-quality generation.
Grouping Factor. Table 4 and Figure 2 demonstrate that our grouping strategy significantly enhances both performance and computational efficiency. Contrary to degrading generation, grouping substantially improves both speech understanding (S2T) and speech-to-speech generation (S2M). For instance, increasing the grouping factor from 1 to 5 raises the S2T score from 55.67 to 63.33 (a 13.7% relative gain). The S2M (T/S) score sees a significant improvement, peaking at 37.67 / 28.00 with a grouping factor of 5. Furthermore, as shown in Figure 2, using a grouping factor of 5 instead of 1 reduces nearly 50% GPU hours in each setting, suggesting the efficiency of grouping machinism.
@@
Speech Refinement. As shown in Table 5, the Speech Refinement Head (SRH) is remarkably effective for speech token generation tasks. By comparing the model with SRH (w/o. SRH-Pretraining) to one without (w/o. SRH), we observe that adding SRH achieves a 76.9% relative improvement in S2M (T) (from 21.67 to 38.33) and a 31.2% relative improvement in T2M (T) (from 45.22 to 59.33). The text-only performance (T2T) remains stable, confirming that SRH enhances speech generation without interfering with text processing pathways. Our dual-resolution architecture effectively combines these components, using grouping for efficient comprehension and SRH at the raw frame resolution for speech generation.
@@
To examine the impact of pretraining, we remove the SRH pretraining step and retrain the model. As shown in Table 5, removing pretraining (comparing w/o. CSE with w/o. SRH-Pretraining) has the most significant impact on speech generation, causing S2M (T) performance to drop by 37.8% and T2M (T) by 15.2%. The effect on S2T is smaller (10.2% drop), and it is negligible for T2T. This underscores the critical importance of SRH pretraining for refining speech token generation.
CoM-Mixing Training Strategy.
@@
As shown in Table 5, where tasks guided by contextual system prompts (STC/SAC/SUC) demonstrate substantial improvements over direct S2M generation. For example, STC (T) achieves a score of 75.67, significantly surpassing the S2M (T) baseline of 68.67. This shows the model successfully learns to adopt the generation paradigm guided by system prompts. The data augmentation effect is confirmed by ablating the CoM-Mixing strategy entirely. Training without it (w/o. CoM-Mixing Training) results in a 15.5% relative performance degradation in S2M (T) (from 68.67 to 58.00), confirming the value of our mixed-modality training approach.
Data Scaling
To investigate the impact of data scaling, we conduct experiments by progressively expanding the S2S training data on DrVoice-Small (w/o. Continous Speech Encoder). Each S2S instance can be augmented into 7 distinct multimodal interaction patterns. As illustrated in Figure 3, expanding the dataset from 3.6K to 17.8K samples yields consistent improvements. The performance of T2M(S) exhibits nearly linear growth, while other patterns, despite showing slower growth rates, still demonstrate upward trends in the figure, suggesting potential for further performance enhancement through additional data.
5 Conclusions and Future Work

### v3

@@ 3.3 Training Strategy
Core-Cocktail Training. We identify a learning rate dilemma: employing a high learning rate significantly compromises the performance of the MLLM, whereas using a low learning rate leads to training stagnation, with the loss decreasing quite slowly. To overcome this optimization challenge, we develop a specialized two-stage training strategy, termed “Core-Cocktail”. The first stage involves full fine-tuning the entire MLLM with a relatively high learning rate. The goal of this phase is to rapidly move the model’s parameters into a more favorable region of the loss landscape. However, to counteract the potential performance degradation of the MLLM caused by this aggressive initial training, an intermediate merging step is introduced. Drawing inspiration from Xiao et al. (2024), the parameters of the MLLM trained in the first stage ($M_{1}$) are merged with the parameters of the base, pre-trained LLM ($M_{0}$). This merging creates a new, interpolated model, $M_{r}$, as defined by the following equation:
$M_{r}\leftarrow\alpha M_{1}+(1-\alpha)M_{0}$
where $M_{r}$ denotes the merged model and $\alpha$ is the interpolation weight. The merging step effectively re-integrates the robust knowledge of the base LLM. A smaller $\alpha$ corresponds to a greater preservation of the base LLM’s capabilities. The second stage conducts full fine-tuning on the merged model $M_{r}$ using a small learning rate. The core-cocktail training approach allows for careful and precise optimization, refining the model’s performance without instability from a high learning rate.
@@ 4 Experiments
Table 2: Performance Comparison on various benchmarks in terms of benchmark-specific metrics (the best result in each row is in bold). With the exception of GLM4-Voice, whose results are cited from Xu et al. (2025) and Chen et al. (2024b), all other results were generated by running inference on the released checkpoints. FR(In/Out) denotes the input speech frame rate and the output speech plus text frame rate for the LLM backbone. $\tau$ denotes the average number of text tokens corresponding to one second of speech.
GLM4-Voice
MiniCPM
-o 2.6
Baichuan
-Omni-1.5
Qwen2.5
-Omni
Kimi
-Audio
Step-Audio2
-Mini
DrVoice
@@
OpenAudioBench (S2T)
AlpacaEval 57.89 64.10 77.90 72.76 75.73 59.60 78.34
Llama Q. 76.00 78.00 78.50 75.33 79.33 75.00 80.33
Reasoning QA 47.43 38.60 50.00 63.76 58.02 46.04 57.92
TriviaQA 51.80 63.00 57.20 57.06 62.10 57.70 61.50
Web Q. 55.40 69.20 59.10 62.80 70.20 65.10 68.10
Overall 57.70 62.58 64.54 66.34 69.08 60.69 69.24
VoiceBench (S2T)
AlpacaEval 3.97 4.42 4.50 4.33 4.46 4.17 4.52
CommonEval 3.42 4.15 4.05 3.84 3.97 3.00 3.77
SD-QA 36.98 50.72 43.40 57.41 63.12 56.06 68.54
MMSU 39.75 54.78 57.25 56.38 62.17 52.18 60.31
OpenBookQA 53.41 78.02 74.51 79.12 83.52 64.18 79.56
IFEval 52.80 49.25 54.54 53.88 61.10 38.01 59.30
AdvBench 88.08 97.69 97.31 99.62 100.00 93.08 98.65
Overall 59.83 71.69 71.14 72.83 76.93 63.84 76.02
UltraEval-Audio (S2S)
AlpacaEval 51.00 51.00 58.69 56.10 44.20 51.72 49.65
Llama Q. 50.00 61.00 67.33 66.30 57.33 67.67 68.00
TriviaQA 36.40 40.20 30.57 40.52 35.71 33.50 35.35
Web Q. 32.00 40.00 38.09 38.93 33.90 34.65 37.65
Overall 42.35 48.05 48.67 50.46 42.79 46.89 47.66
Big Bench Audio (S2T & S2S)
S2T 44.8 56.2 47.1 54.2 59.4 50.9 71.6
S2S 42.7 55.4 44.6 53.6 51.0 47.5 60.9
Overall 43.8 55.8 45.8 53.9 55.2 49.2 66.3
Table 3: Speech quality performance on the collections of UltraEval-Audio, in terms of UTMOS for assessing overall speech quality and ASR-WER for assessing alignment between generated speech and text. FR(In/Out) indicates the input speech frame rate and the output (speech + text) frame rate for the LLM backbone, while $\tau$ represents the average text tokens per second of speech.
@@
DrVoice 5/5 4.29 11.2
@@ 4.1 Experimental Setup
Datasets. Approximately 100K hours of audio-text paired data for speech-text modality alignment is adopted for pre-training Speech Refined Head. For DrVoice post-training, we first synthesize speech for about 3B text tokens using CosyVoice (Du et al., 2024a, b, 2025), then select about 26K hours for speech-to-speech conversation and about 20K hours user speech plus 1.3B assistant tokens for speech-to-text conversation, based on Word Error Rate (WER) of the synthesized speech. Furthermore, to enhance the model’s comprehension of real-world speech, the training data was augmented with about 10K hours of English Automatic Speech Recognition (ASR) data, comprising several corpora, including Common Voice (Ardila et al., 2020), MELD (Poria et al., 2019), LibriSpeech (Panayotov et al., 2015), SPGISpeech (O’Neill et al., 2021), and Voxpopuli (Wang et al., 2021). Following prior works (Yao et al., 2024; KimiTeam et al., 2025; OpenAI, 2024b), we evaluate the performance on the widely used benchmarks, VoiceBench (Chen et al., 2024b) and OpenAudioBench  [^1] https://huggingface.co/datasets/baichuan-inc/OpenAudioBench for Speech-To-Text ($S\rightarrow T$) evaluation, UltraEval-Audio  [^2] https://github.com/OpenBMB/UltraEval-Audio for Speech-to-Speech ($S\rightarrow S$) evaluation, and Big Bench Audio  [^3] https://huggingface.co/datasets/ArtificialAnalysis/big_bench_audio for both evaluations.
@@
Evaluation Metrics. Evaluations adhere to the established protocols for each respective benchmark. Specifically, for the open-ended QA tasks on AlpacaEval and CommonEval, G-Eval (Liu et al., 2023) is used for scoring. For AdvBench, the Refusal Rate is reported, while performance on all other benchmarks is assessed with Accuracy. The generated speech is transcribed using Whisper-v3-large model (Radford et al., 2022), then WER (denoted by ASR-WER) of transcripts is computed against the generated text to assess the alignment between generated speech and text. UTMOS (Saeki et al., 2022) is used to evaluate the overall speech quality, following Zeng et al. (2025).
@@
Baselines. Representive and competitive open-source audio language models are selected as baselines to cover diverse modeling paradigms: Text-Driven models including MiniCPM-o 2.6 (8B) (Yao et al., 2024) and Qwen2.5-Omni (7B) (Xu et al., 2025). Among Joint Speech-Text models, interleaved models including GLM-4-Voice (9B) (Zeng et al., 2025), Baichuan-Omni-1.5 (7B) (Li et al., 2025), and Step-Audio2-Mini (8B) (Wu et al., 2025), alongside the parallel model Kimi-Audio (7B) (KimiTeam et al., 2025). This suite enables systematic comparisons across mainstream speech-text modeling strategies.
Implementation Details can be found in Appendix A.
4.2 Main Results
@@
Computational Efficiency and Speech Quality. A key innovation of DrVoice is its remarkable computational efficiency, highlighted in Table 2 and Table 3. As shown in the FR (In/Out) rows, DrVoice operates at a frame rate of 5/5, indicating that the LLM backbone processes only 5 audio tokens per second for both input and output. This is a substantial reduction compared to other models, which operate at frame rates of 12.5 or 25, thereby significantly lowering computational requirements and potential latency. Crucially, this efficiency does not come at the cost of speech quality. Table 3 shows that despite its low frame rate, DrVoice produces high-quality and well-aligned speech. It achieves UTMOS score of 4.29 for overall speech quality, which is competitive with top-performing models like Qwen2.5-Omni (4.28) and superior to others like Kimi-Audio (3.06). With an ASR-WER of 11.2, the model shows robust speech-text alignment and surpasses several baselines. Still, it underperforms compared to Qwen2.5-Omni. A potential explanation lies in the architectural design: Qwen2.5-Omni’s feeding text directly to its “Talker” module, whereas the proposed model only sends hidden states to SRH. Therefore, introducing text to the speech refined head is a logical next step for improving textual control and lowering the ASR-WER. This unique combination of SOTA performance, high-quality speech generation, and unparalleled efficiency makes DrVoice a highly powerful and practical model for real-world applications.
4.3 Ablation and Analyses
We conduct extensive ablation study and analyses. For computational efficiency, some experiments are performed on DrVoice-Small (1.5B). More analyses about the data quality and data scaling can be found in Appendix C.
Core-Cocktail Training Strategy. While Stage 1 training causes a performance drop from text baseline of 81.77 to 70.19, Stage 2 reverses this trend, recovering the performance to 74.73. This result confirms that the strategy effectively counteracts the initial degradation and leads to an optimized model. Further details are available in Appendix C.
@@ 4.3 Ablation and Analyses
Continuous Speech Encoder. As shown in Table 4, the Continuous Speech Encoder (CSE) is vital for tasks involving speech inputs. Removing it from DrVoice-Small leads to a significant performance drop in speech understanding and speech generation. Specifically, the S2T score decreases from 72.33 to 62.33 (13.8% relatively), and the S2M (T) score falls from 68.67 to 61.67 (10.2% relatively). In contrast, the impact on text-only tasks is minimal, with the T2T score only slightly decreasing from 75.33 to 74.00. This confirms the CSE’s effectiveness in representing speech.
@@
Dual-Resolution Speech Representations (DRSR). Our dual-resolution approach combines input grouping for efficiency and comprehension with a refinement head for high-quality generation. 1) Grouping Factor. Contrary to degrading generation, grouping substantially improves both speech understanding (S2T) and speech-to-speech generation (S2M). For instance, increasing the grouping factor from 1 to 5 raises the S2T score from 55.67 to 63.33 (a 13.7% relative gain). The S2M (T/S) score sees a significant improvement, peaking at 37.67 / 28.00 with a grouping factor of 5. Furthermore, using a grouping factor of 5 instead of 1 reduces nearly 50% GPU hours in each setting, proving the efficiency of grouping machinism. Detail results can be found in Appendix C. 2) Speech Refinement Head. As shown in Table 4, the Speech Refinement Head (SRH) is remarkably effective for speech token generation tasks. By comparing the model with SRH (w/o. SRH-Pretraining) to one without (w/o. SRH), we observe that adding SRH achieves a 76.9% relative improvement in S2M (T) (from 21.67 to 38.33) and a 31.2% relative improvement in T2M (T) (from 45.22 to 59.33). The text-only performance (T2T) remains stable, confirming that SRH enhances speech generation without interfering with text processing pathways. Our dual-resolution architecture effectively combines these components, using grouping for efficient comprehension and SRH at the raw frame resolution for speech generation.
@@
SRH-Pretraining. To examine the impact of pretraining, we remove the SRH pretraining step and retrain the model. As shown in Table 4, removing pretraining (comparing w/o. CSE with w/o. SRH-Pretraining) has the most significant impact on speech generation, causing S2M (T) performance to drop by 37.8% and T2M (T) by 15.2%. The effect on S2T is smaller (10.2% drop), and negligible for T2T. This underscores the critical importance of SRH pretraining for refining speech token generation.
@@
CoM-Mixing Training Strategy. As shown in Table 4, where tasks guided by contextual system prompts (STC/SAC/SUC) demonstrate substantial improvements over direct S2M generation. For example, STC (T) achieves a score of 75.67, significantly surpassing the S2M (T) baseline of 68.67. This shows the model successfully learns to adopt the generation paradigm guided by system prompts. The data augmentation effect is confirmed by ablating the CoM-Mixing strategy entirely. Training without it (w/o. CoM-Mixing Training) results in a 15.5% relative performance degradation in S2M (T) (from 68.67 to 58.00), confirming the value of our mixed-modality training approach.
5 Conclusions